OpenStudy (anonymous):

Which of the following is a irrational number and why? A.1/32 B.3.141592 C.check mark 16 D.-22 Algebra Help!

OpenStudy (anonymous):

lol its C.

OpenStudy (anonymous):

checkmark 16

OpenStudy (anonymous):

i belive so

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Why

OpenStudy (agent0smith):

wtf is check mark 16. Square root of 16? That's 4. Is 4 irrational? 3.141592... assuming this is pi, is irrational.

OpenStudy (anonymous):

because it is not a real #

OpenStudy (agent0smith):

^ yes the square of 16 is... Irrational means it's a never-ending decimal that never repeats.

OpenStudy (anonymous):

SO what is it Agent0smith

OpenStudy (agent0smith):

I already said what the answer is above...

OpenStudy (anonymous):

Pi?

OpenStudy (anonymous):

pi is irrational or3.141592.....is irrational. but 3.141592 is rational.

OpenStudy (anonymous):

so what is the answer surjeithayer

OpenStudy (anonymous):

i don't know what is meant by check 16.

OpenStudy (ybarrap):

Any number that can not be represented by a/b where a and b are integers is irrational. Can \(\pi\) be represented by as fraction a/b? Not approximately, exactly. Numbers that can be represented as a/b are rational. Check 16 sounds like you are interpreting your options as \(\sqrt{}16=\sqrt{16}\).

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

that means there is no number . If it is 3.141592..... never ending then only it is irrational .

OpenStudy (ybarrap):

can PI (\(\pi=3.141592653589793238462643383279502884\\1971693993751058209749445923078164062862089986\\2803482534211706798214808651328230664709384460\\95505822317253594081284811174502841027019385211\\0555964462294895493038196442881...)=\cfrac{a}{b}\) for some a and b? The digits of PI goes on forever and there is no a and b that can represent such a number

OpenStudy (agent0smith):

lol @boyeatyourCHICKEN the answer got stated many times. By me and others.